# Community detection in network using Szegedy quantum walk

### v1

@@ 3 Community detection on graphs
[…]
$\overrightarrow{P}_{t}=(p_{t}(\overrightarrow{e_{1}}),p_{t}(\overrightarrow{e_{2}}),\dots,p_{t}(\overrightarrow{e_{n}}),p_{t}(\overrightarrow{e_{1}^{-1}}),p_{t}(\overrightarrow{e_{2}^{-1}}),\dots,p_{t}(\overrightarrow{e_{m}^{-1}})).$ (9)
Note that, every edge $e$ of the undirected graph $G$ contributed two directed edges $\overrightarrow{e}$ and $\overrightarrow{e^{-1}}$, in the directed graph $\overrightarrow{G}$. Thus, we can determine the edge probability $e$ in $G$ at $t$-th step of quantum walk as $p_{t}(e)=p_{t}(\overrightarrow{e})+p_{t}(\overrightarrow{e^{-1}})$. Therefore the edge probability vector of the undirected graph $G$ at the $t$-th step of quantum walk is represented by
@@
where $p_{t}(e_{i})=p_{t}(\overrightarrow{e_{i}})+p_{t}(\overrightarrow{e_{i}^{-1}})$. Note that, due to the unitary dynamics, $P_{t}$ need not converge to a limiting distribution for $t\rightarrow\infty$ [36]. But, the time average probability distribution converges. Hence, we have the following definitions.
@@
The edge probability vector at time $T$ of a graph $G$ is defined by
$\begin{split}\Pi^{(T)}&=\left(\pi_{1}^{(T)},\pi_{2}^{(T)},\dots\pi_{m}^{(T)}\right)\\ &=\frac{1}{(T+1)}\sum_{t=0}^{T}P_{t}=\frac{1}{(T+1)}\left(\sum_{t=0}^{T}p_{t}(e_{1}),\sum_{t=0}^{T}p_{t}(e_{2}),\dots,\sum_{t=0}^{T}p_{t}(e_{m})\right),\end{split}$
where $\pi_{k}^{(T)}=\frac{1}{(T+1)}\sum_{t=0}^{T}p_{t}(e_{k})$ for $k=1,2,\dots m$.
Note that, the edge probability vector at time $T$ changes over time. It proceeds to a limiting distribution when $T\rightarrow\infty$. Hence, we can define the edge probability vector of a graph.
@@
The edge-probability vector of a graph $G$ is defined by
$\Pi=(p(e_{1}),p(e_{2}),\dots,p(e_{m}))=\lim_{T\to\infty}\frac{1}{(T+1)}\sum_{t=0}^{T}P_{t}=\lim_{T\to\infty}\frac{1}{(T+1)}\left(\sum_{t=0}^{T}p_{t}(e_{1}),\sum_{t=0}^{T}p_{t}(e_{2}),\dots,\sum_{t=0}^{T}p_{t}(e_{m})\right).$
For our computation, rather than calculating $T\to\infty$, we find $\frac{1}{T+1}\sum_{t=0}^{T}P_{t}$ for a finite $T$ to construct the limiting probability distribution. We apply the standard Euclidean distance to differentiate two edges probability vectors $\Pi^{(T)}$ and $\Pi^{(T-1)}$, which is determined by
$\begin{split}\left\lVert\Pi^{(T)}-\Pi^{(T-1)}\right\rVert&=\left\lVert\frac{1}{T+1}\sum_{t=0}^{T}P_{t}-\frac{1}{T}\sum_{t=0}^{T-1}P_{t}\right\rVert\\[6.0pt] &=\sqrt{(\pi_{1}^{(T)}-\pi_{1}^{(T-1)})^{2}+(\pi_{2}^{(T)}-\pi_{2}^{(T-1)})^{2}+\cdots+(\pi_{m}^{(T)}-\pi_{m}^{(T-1)})^{2}}.\end{split}$ (11)
To construct the edge probability vector of a graph $G$, we compute a sequence of edge probability vectors $\{\Pi^{(t)}\}_{t=0}^{T}$ for time $t$, where
@@
Here, we consider $\epsilon$ as a numerical error which is usually a very small number. For our numerical experiments, we consider $\epsilon=0.0001$. The consideration of numerical error assists us to find the accurate value of $\Pi$ to three decimal places. In 1(b), we present the edge-probabilities of the graph in 1(a). The convergence of the probability vectors are discussed in Appendix B.
@@
Now, we include a number graph theoretic ideas in construction of our communities. A vertex with degree $1$ is adjacent to only one vertex, which may belong to a community. A vertex with degree two lies on a path between two neighboring vertices. These three vertices may form a single community, or the neighboring vertices may belong to two different communities. But, a vertex with greater degree is adjacent to many other vertices. Thus, it belong to only one community. It insists us to arrange the vertices in a descending order of their degrees. In every step, we consider the highest degree vertex among the vertices which are not allotted to a community and generate a community for it.
In addition, the above paragraph justifies that the community containing a vertex with highest degree also contains all its neighboring vertices. But, there may be a number of vertices in a community, which are not an immediate neighbor of the vertex with highest degree in it. Also, there are non-neighboring vertices which are not in the community. When a vertex is in the community we observe that most of its neighbours are also in the same community. Therefore, we need a condition for including a given vertex in a particular community. Recall from subsection 2.1 that $N(v_{i})=\{v:(u,v)\in E(G)\}$. Let $\operatorname{Nbd}(v_{i})$ be the neighborhood of $v_{i}$ in the community of vertices containing $v_{i}$. If $|N(v_{j})\cap\operatorname{Nbd}(v_{i})|\geq\frac{|N(v_{j})|}{2}$ holds, we assume that $v_{j}$ is in the community of $v_{i}$.
We make the final selection of a vertex in a community based on the edge probability vector $\Pi$ generated by quantum walk. We determine the shortest path between $v_{i}$ and $v_{j}$. A path consists of a number of incident edges. The walker follows the path means it follows all the subsequent edges. As there is no non-Markovian componenet in our quantum walk, the probability of getting the walker on the edges of a path is independent to each other. Therefore, the probability of getting the walker on a path is the product of the probability of getting the walker on the edges. It leads us to define the path weight between two vertices.
@@
Note that, as all $0<p(e_{i})<1$ and $d_{G}(v)\geq 1$ we have $0<PW(u,v)<1$ for all pairs of vertices $u$ and $v$. The explanation for dividing $\frac{1}{d_{G}(v)}$ will be discussed after Procedure Procedure 1.
Now, we are in a position to determine the communities in a network. Given a vertex $v_{i}$ the community consists of $v_{i}$ is denoted by $C(v_{i})$. Initially, $C(v_{i})=\{v_{i}\}$. Then, $C(v_{i})$ grows maintaining some conditions. We choose the vertices $v_{i}$ according to their degrees. Our method is described in the procedure mentioned below.
@@
Recall that $N(v)$ denotes the subset of vertices which are adjacent to $v$ in $G$. The procedure for detecting the communities consists of the following steps:
@@
    Arrange the vertices of the graph $G$ in the descending order of the vertex degrees. Therefore, the vertex set of $G$ is $V=\{v_{1},v_{2},\dots v_{n}\}$, where $d_{G}(v_{i})\geq d_{G}(v_{i+1})$ for $i=1,2,\dots(n-1)$.
@@
    Repeat the following steps if $V\neq\emptyset$.
@@
    Return the communities of $C(v_{i})$ which are formed in Step 2.
It is clear from the Procedure 1 mentioned above that we construct a set of vertices $\operatorname{Nbd}(v_{i})$ based on equation (14), which assists us to construct the community $C(v_{i})$. The quantum walk is useful to calculate the probability values in finding $PW(v_{i},v_{j})$ mentioned in equation (13).
In the above procedure $q$ is a parameter. If $q=0$ then the number of communities is equal to the number of vertices in the graph, that is every community consists of single vertices. When $q>\max_{(u,v)\in E(G)}\{PW(u,v)\}$ then $C(v_{i})=\operatorname{Nbd}(v_{i})$.
In equation (13), we multiply $\frac{1}{d_{G}(v)}$ to normalize node influence. It provides more preference to a high-degree node to be added in $C(v_{i})$, because a node with higher degree reduces the product $\prod_{i=1}^{n}p(e_{i})$ than a vertex with lower degree.
@@
    For any vertex $v_{j}\in C(v_{i})$ if $d_{C(v_{i})}(v_{j})\leq\frac{d_{G}(v_{j})}{2}$ then remove $v_{j}$ from $C(v_{i})$. Here, $C(v_{i})$ denotes the induced subgraph of $G$, generated by $C(v_{i})$.
@@
Let $G$ be a simple undirected graph with $n$ vertices and $m$ edges. To find the communities in $G$ perform the following steps:
@@
    Generation of a directed graph: Convert the graph $G$ to a directed graph $\overrightarrow{G}$ by assigning opposite orientations on every undirected edge.
@@
    Generation of limiting probability distribution: For a fixed error term $\epsilon$, we apply equation (12) to stop quantum walk and generate the limiting probability distribution $\Pi$ on the edges mentioned in Definition 2.
@@ 4 Numerical experiments
We have already mentioned that not all graphs admits community structure in it. We consider a number of graphs with a community structure. We also consider a graph having no trivial community structure. We initiate our discussion with a few standard examples. Some of the graphs considered in this article are the random graphs. Due to their generation procedure, they may have directed edges or loops. But, our procedure is applicable for undirected simple graphs. During the network generation process, we remove the loops and directed edges if exist.
For every graph we include a figure with at least four sub-figures. Below, we describe them:
  1. 1.
    In Sub-figure (a), we draw the example graph, which we consider as our input.
  2. 2.
    In Sub-figure (b), we include a bar diagram representing the limiting edge-probability distribution $\Pi$, mentioned in Definition 2 after completion of quantum walk.
  3. 3.
    In Sub-figure (c), we draw a bar diagram mentioning the path weights between selected vertices in the $\operatorname{Nbd}$ of the community generating vertex. The equation of path probability is mentioned in equation (13).
  4. 4.
    In Sub-figure (d), we draw the graph with communities in it. The vertices in same community have same color. There are communities with single vertices. We mark all of them in gray.
We use NetworkX, a Python library, for all the calculations related to graphs [37].
@@ 4.1 Barbell graphs
A barbell graph [38] is generated by joining $r$ copies of a complete graph $K_{s}$ with a particular vertex, called central vertex. We join the central vertex to the $r$ vertices belonging to different $K_{s}$ by $r$ edges. For example in 2(a), we depict a barbell graph, that has two complete graphs each of five vertices. We add two vertices from each of the $K_{5}$ graphs with a vertex $5$ by two edges. Here vertex $5$ acts as the central vertex. Therefore, the barbell graph under our consideration has eleven vertices and twenty-two edges.
Considering $q=\frac{1}{m}=0.0455$ in Procedure 3, we obtain three distinct communities in the barbel graph network. Two distinct communities arise from two complete graphs.The central vertex forms a community with a single vertex. The coverage values corresponding to $q=0.0455$ is $0.909$. In 2(d) we present the communities generated by Procedure 3.
@@
(b) Edge probabilities in the limiting distribution $\Pi$ are depicted as bar diagram for the graph in Subfigure 2(a).
(c) Path weight between different vertices of the for generating the communities from the graph in 2(a).
@@ 4.2 Relaxed caveman graph
The idea of caveman graphs arises from the social networks [39, 40]. The connected caveman graph is generated by creating $k$ complete subgraphs with equal number of vertices in each. Then a single edge in each clique is rewired to a node in an adjacent clique. These cliques are called caves. The relaxed caveman graph is a variant of the connected caveman graphs. The network is composed of $k$ nodes organized into $k$ caves. For every edge there is a probability $p$ of rewiring, that is removing an edge and adding it to any vertex keeping one end point fixed. A relaxed caveman graph is depicted in 3(a).
The example graph, depicted in 3(a), has $15$ vertices and $30$ edges. The vertices are distributed into $3$ caves. Every cave contains $5$ vertices in it. We consider that the rewiring probability is $0.2$. The example graph has $4$ inter-community edges.
Note that, in this case $\frac{1}{m}=0.033$. The communities are marked with three different colors in Sub-figure 3. Decreasing the value of $q_{2}$ to $0.0166$ we find more communities in our network. Coverage are $0.866$ and $0.766$ for $q=0.0333$ and $0.0166$ respectively.
@@ 4.3 Planted $l$-partition graph
The planted partition models of networks offer another controllable artificial input for evaluating and benchmark the community detection algorithms. The Erdős-Rényi model of random graphs [41] with $n$ vertices is generated by choosing each of the $\binom{n}{2}$ possible edges independently and randomly to be included in the graph with probability $p$. The planted $l$-partition model is a modified version of the Erdős–Rényi model, which is specifically designed to generate communities in graph. To generate $l$ communities in a graph with $k$ vertices in each, we need $n=l\times k$ vertices. Now, we fix two probabilities $p$ and $q$ with $p>q$. Existence of an edge between two vertices belonging to same cluster is determined by probability $p$. An edge between two vertices from two different cluster occurs in probability $q$. A planted $l$-partition graph is depicted in 4(a).
We consider a sample of a planted partition graph, which in 4(a). It has twelve vertices and fourteen edges. Also, values of $p_{1}$ and $p_{2}$ are $0.4$ and $0.1$, respectively. The vertices are divided into three subset each having four vertices in it.
The vertices with the highest degrees in the graph are identified as 0, and $7$. Among them, vertex 0 is randomly selected as the first vertex to initiate Procedure 1 for constructing a community. A bar diagram is employed to illustrate the route probability distribution. Based on the graphs structural connectivity, a distinct community can be obtained for $q>0.071$. The coverage values corresponding to $q=0.071$ is $0.857$.
@@
(c) Path weight between different vertices of the for generating the communities from the graph in 4(a).
@@ 4.4 Zachary’s karate club graph
The Zachary’s karate-club network [42] is another well-studied graph considered in the literature of community detection [33, 32]. This is a social network of university karate club members. Each of the $34$ vertices represents a club member. There are $78$ edges represent the relations between the members. Article [33] presents three communities in the Zachary’s karate-club network. In contrast, article [32] utilize a different community detection technique to analyze the same network that found two communities only.
Considering $q=\frac{1}{m}=0.0128$ in Procedure 3, we obtain three distinct communities in the Zachary’s karate-club network. Two communities have fourteen and eighteen vertices, and one community consists of a single vertex. The coverage values corresponding to $q=0.0128$ is found to be $0.859$. In 5(d) we present the communities determined by Procedure 3. In contrast, the article [33] determines three communities having five, ten, and nineteen vertices. Furthermore, article [32] develops two communities consist of sixteen and eighteen vertices.
@@
(b) The bar diagram for the graph in Subfigure 5(a) represents the edge probability in the limiting distribution $\Pi$.
(c) In 5(a), the path weight between various vertices of the graph is used to generate the communities.
(d) Three distinct colors are used to extract communities from the graph in Subfigure 5(a).

### v2

@@ 3 Community detection on graphs
Using equation (6), we construct an edge probability vector for $\overrightarrow{G}$ at the $t$-th step of quantum walk as
$\overrightarrow{P}_{t}$ $=\bigl(p_{t}(\overrightarrow{e}_{1}),\ldots,p_{t}(\overrightarrow{e}_{m}),p_{t}(\overrightarrow{e}_{1}^{-1}),\ldots,p_{t}(\overrightarrow{e}_{m}^{-1})\bigr).$ (7)
Note that, every edge $e$ in $G$ contributes two directed edges $\overrightarrow{e}$ and $\overrightarrow{e^{-1}}$, in $\overrightarrow{G}$. Thus, we determine the probability of $e$ in $G$ at time $t$ as $p_{t}(e)=p_{t}(\overrightarrow{e})+p_{t}(\overrightarrow{e^{-1}})$. Therefore, the edge probability vector of $G$ at time $t$ as
@@
where $p_{t}(e_{i})=p_{t}(\overrightarrow{e_{i}})+p_{t}(\overrightarrow{e_{i}^{-1}})$. Note that, due to the unitary dynamics, $P_{t}$ need not converge to a limiting distribution as $t\rightarrow\infty$ [42]. But, the time average probability distribution converges. Hence, we have the following definitions.
@@
The edge-probability vector at time $T$ of a graph $G$ is defined by
$\begin{split}&\Pi^{(T)}=\bigl(\pi_{1}^{(T)},\pi_{2}^{(T)},\ldots,\pi_{m}^{(T)}\bigr)=\frac{1}{T+1}\sum_{t=0}^{T}P_{t}\\ &=\frac{1}{T+1}\bigl(\sum_{t=0}^{T}p_{t}(e_{1}),\sum_{t=0}^{T}p_{t}(e_{2}),\ldots,\sum_{t=0}^{T}p_{t}(e_{m})\bigr),\end{split}$
where $\pi_{k}^{(T)}=\frac{1}{T+1}\sum_{t=0}^{T}p_{t}(e_{k})$ for $k=1,2,\ldots,m$.
Note that, the edge probability vector proceeds to a limiting distribution when $T\rightarrow\infty$, which we define as the edge probability vector of a graph.
@@
For a graph $G$ the edge-probability vector
$\begin{split}&\Pi=\bigl(p(e_{1}),p(e_{2}),\ldots,p(e_{m})\bigr)=\lim_{T\to\infty}\frac{1}{T+1}\sum_{t=0}^{T}P_{t}\\ &=\lim_{T\to\infty}\frac{1}{T+1}\Bigl(\sum_{t=0}^{T}p_{t}(e_{1}),\sum_{t=0}^{T}p_{t}(e_{2}),\ldots,\sum_{t=0}^{T}p_{t}(e_{m})\Bigr).\end{split}$
For our computation, rather than calculating $T\to\infty$, we find $\frac{1}{T+1}\sum_{t=0}^{T}P_{t}$ for a finite $T$. We apply the standard Euclidean distance to differentiate two edge probability vectors $\Pi^{(T)}$ and $\Pi^{(T-1)}$, which is
$\left\lVert\Pi^{(T)}-\Pi^{(T-1)}\right\rVert$ $=\left\lVert\frac{1}{T+1}\sum_{t=0}^{T}P_{t}-\frac{1}{T}\sum_{t=0}^{T-1}P_{t}\right\rVert.$ (9)
To work out the edge probability vector of $G$, we compute a sequence of edge probability vectors $\{\Pi^{(t)}\}_{t=0}^{T}$, where
@@
In our numerical experiments, $\epsilon=0.0001$, which assists us to find the accurate value of $\Pi$ to three decimal places. In 1(b), we present the edge-probabilities of the graph in 1(a). The convergence of the probability vectors are discussed in Appendix.
@@
Now, we include a few graph-theoretic ideas. A vertex with degree $1$ is adjacent to only one vertex, which may belong to a community. A vertex with degree two lies on a path between two neighboring vertices. These three vertices may form a single community, or the neighboring vertices may belong to two different communities. But, a vertex with greater degree is adjacent to many other vertices. Thus, it belongs to only one community. It insists us to arrange the vertices in a descending order of their degrees. In every step, we consider the highest degree vertex among the vertices which are not allotted to a community.
There may be a number of vertices in a community, which are not an immediate neighbor of the vertex with highest degree in it. Also, there are non-neighboring vertices which are not in the community. When a vertex is in the community we observe that most of its neighbours are also in the same community. Therefore, we need a condition for including a given vertex in a particular community. Recall that $N(v_{i})=\{v:(u,v)\in E(G)\}$. Let $\operatorname{Nbd}(v_{i})$ be the neighborhood of $v_{i}$ in the community of vertices containing $v_{i}$. If $|N(v_{j})\cap\operatorname{Nbd}(v_{i})|\geq\frac{|N(v_{j})|}{2}$ holds, we assume that $v_{j}$ is in the community of $v_{i}$.
The final selection of a vertex in a community depends on the edge probability vector $\Pi$. We determine the shortest path between $v_{i}$ and $v_{j}$. The walker follows the path means it follows all the consecutive edges. As there is no non-Markovian componenet in our quantum walk, the probability of getting the walker on the edges of a path is independent to each other. Therefore, the probability of getting the walker on a path is the product of the probabilities. It leads us to define the path weight between two vertices.
@@
Let $v_{i}$ be a vertex of maximum degree among the vertices which are not included in a community. The community consists of $v_{i}$ is denoted by $C(v_{i})$. Initially, $C(v_{i})=\{v_{i}\}$. The Procedure mentioned below add the neighboring vertices of $v_{i}$ in $C(v_{i})$.
@@
Follow the steps mentioned below:
@@
    Arrange the vertices of $G$ as $V=\{v_{1},v_{2},\dots v_{n}\}$, where $d_{G}(v_{i})\geq d_{G}(v_{i+1})$.
@@
    Repeat the following steps if $V\neq\emptyset$:
@@
    Return the communities of $C(v_{i})$.
In Procedure 1 the quantum walk is useful to calculate the probability values in finding $PW(v_{i},v_{j})$ mentioned in Definition 3. Also, $q$ is a parameter. If $q=0$ then the number of communities is equal to the number of vertices in the graph. When $q>\max_{(u,v)\in E(G)}\{PW(u,v)\}$ then $C(v_{i})=\operatorname{Nbd}(v_{i})$.
In Definition 3, we multiply $\frac{1}{d_{G}(v)}$ to normalize node influence. It provides more preference to a high-degree node to be added in $C(v_{i})$, because a node with higher degree reduces the product $\prod_{i=1}^{n}p(e_{i})$ than a vertex with lower degree.
@@
    For any vertex $v_{j}\in C(v_{i})$ if $d_{C(v_{i})}(v_{j})\leq\frac{d_{G}(v_{j})}{2}$ then remove $v_{j}$ from $C(v_{i})$.
@@
To find the communities in a simple graph $G$ perform the following steps:
@@
    Generation of a directed graph: Generate a directed graph $\overrightarrow{G}$ from $G$ by assigning opposite orientations on every undirected edge.
@@
    Generation of limiting probability distribution: For a fixed error term $\epsilon$, we apply equation (10) to stop quantum walk and generate the limiting probability distribution $\Pi$ mentioned in Definition 2.
@@ 4 Numerical experiments
We consider a number of standard examples of graphs with a community structure, which are considered in literature. Some of them are the random graphs, which may have directed edges or loops. But, our procedure is applicable for undirected simple graphs. Hence, we remove the loops and directed edges if exist. We use NetworkX, a Python library, for all the computations related to graphs [43].
@@ 4.1 Barbell graphs
In 2(a), we consider a barbell graph [44] which is generated by joining two copies of the complete graph $K_{5}$ with a central vertex by two edges. This graph has eleven vertices and twenty-two edges.
Considering $q=\frac{1}{m}=0.0455$ in Procedure 3, we obtain three distinct communities in the barbel graph. Two distinct communities arise from two complete graphs. The central vertex forms a community with a single vertex. The coverage values corresponding to $q=0.0455$ is $0.909$. In 2(c), we present the communities generated by Procedure 3.
@@
(b) Path weight between different vertices of the for generating the communities from the graph in 2(a).
(c) Communities in the graph of Subfigure 2(a) are extracted with three different colors.
@@ 4.2 Relaxed caveman graph
The connected caveman graph [45, 46] is generated by joining complete subgraphs each with $k$ vertices. An edge in a clique is rewired to a node in an adjacent clique with probability $p$. In 3(a) we depict a relaxed caveman graph having $30$ edges, and $15$ vertices which are equally distributed in $3$ caves. We assume $p=0.2$ which creates $4$ inter-community edges.
Putting $q=\frac{1}{m}=0.033$, we find three communities which are presented in Sub-figure 3. Decreasing the value of $q_{2}$ to $0.0166$ we find more communities in our network. Coverage are $0.866$ and $0.766$ for $q=0.0333$ and $0.0166$, respectively.
@@ 4.3 Planted $l$-partition graph
The planted $l$-partition graph [5, 47] has $n=l\times k$ vertices distributed into $l$ communities with $k$ vertices in each. Let $p$ and $q$ with $p>q$ be the probability of existence of the intra-community and inter-community edges, respectively. 4(a) contains a planted $l$-partition graph with $14$ edges and $12$ vertices which are distributed into $3$ communities each having $4$ vertices in it. Also, $p_{1}=0.4$ and $p_{2}=0.1$, respectively.
One of the highest degree vertex 0 is randomly selected as the first vertex to initiate Procedure 1. Considering $q>0.071$, we obtain three communities, where a community contain a single vertex. The coverage values corresponding to $q=0.071$ is $0.857$.
@@
(b) Path weight between different vertices of the for generating the communities from the graph in 4(a).
(c) Communities in the graph of Subfigure 4(a) are extracted with three different colors.
@@ 4.4 Zachary’s karate club graph
The Zachary’s karate-club network [48] is another well-studied graph considered in the literature of community detection [40, 39]. It has $34$ vertices and $78$ edges. Article [40] presents three communities in it. In contrast, article [39] consider the same network that found two communities only.
Considering $q=\frac{1}{m}=0.0128$ in Procedure 3, we obtain three communities in this network. Two communities have fourteen and eighteen vertices, and one community consists of a single vertex. The coverage values corresponding to $q=0.0128$ is $0.859$. In 5(c), we present the communities determined by Procedure 3.
@@
(b) In 5(a), the path weight between various vertices of the graph is used to generate the communities.
(c) Three distinct colors are used to extract communities from the graph in Subfigure 5(a).